OpenStudy (anonymous):

Find the slope of the tangent line to the curve y = cos(2x) when x = pi/6?

hartnn (hartnn):

are u stuck ? u know how to start ?

OpenStudy (anonymous):

I found the derivative to by y' = -2sin2x, but I don't think that's right.

hartnn (hartnn):

that is actually correct.

hartnn (hartnn):

now just put x = pi/6 and u get the slope .

OpenStudy (anonymous):

y' = -2sin(2x) y' = -2sin(2pi/6) y' = -2sin(pi/3) y' = -2(0) y' = 0 ?

hartnn (hartnn):

sin pi/3 = ? its not 0.

OpenStudy (anonymous):

Oh my bad, sin pi/3 = sqrt(3)/2

OpenStudy (anonymous):

So -sqrt(3) since the 2s cancel out?

hartnn (hartnn):

that is correct!

OpenStudy (anonymous):

Thank you very much for taking my through the process!

hartnn (hartnn):

welcome ^_^